From where does this equation: h^2-ab comes in conic sections and what does it represents?

Grade:12

1 Answers

Arun
5 years ago
Β 
Β 
The equation ax2Β + 2hxy + by2Β + 2gx + 2fy + c = 0. Represents a second degree equation where a, h, b doesn’t variables simultaneously.
Β 
Let a β‰  0.
Β 
Now, the above equation becomes
Β 
Β Β Β Β Β Β Β  a2Β x2Β + 2ax (hy + g) = aby2 – 2afy – ac
Β 
on completing the square on the left side, we get,
Β 
Β Β Β Β Β Β Β  a2Β x2Β + 2ax (hy + g) = y2Β (h2 – ab) + 2y (gh – af) + g2 – ac.
Β 
i.e.Β Β Β  (ax + hy + g) =Β + √y2(h2–ab)+2y(gh–af) +(g2–ac)
Β 
We cannot obtain x in terms of y, involving only terms of the first degree, unless the quantity under the radical sign be a perfect square. The condition for this is,
Β 
(gh – af)2Β = (h2 – ab) (g2 – ac)
Β 
hope it clears.
Β 
Now if Delta is 0
Β 
  1. h^2 – ab > 0 intersecting real lines
  2. h^2 – ab = 0. parallel lines
  3. h^2 – ab